I'm confused on how to solve this.
I know i can add it in a calculator bat it would be so long.
I don't know the formula.
Pls. help.
Find the sum: 1 + 2 + 3 ...+ 99 + 100.

Answer by jim_thompson5910(35256)

Use the formula



In your case,




So


Note: you can verify this answer by adding up each number in the sequence (if you're bored), OR you can notice that 1+100=101, 2+99=101, 3+98=101, etc... and notice that there are 50 of these pairings (half of 100). This means that 50*101=5050.